a bathtub contains 70 kg of water at 26C. Ten kg of water at 90C is poured in. what is the final temperature of the mixture.


1
Expert's answer
2021-05-17T12:38:42-0400

"cm_1(t-26)=cm_2(90-t)"


"4218\cdot 70\cdot(t-26)=4218\cdot 10\cdot(90-t)\to"


"t=34\ (°C)" . Answer